OpenStudy (anonymous):

choose the negation of this statement x+y 10 pleaseeee i really need help!

OpenStudy (anonymous):

The sum of x and y is not 10.

OpenStudy (anonymous):

\[\large x+y=10\] ?

OpenStudy (anonymous):

is that what you meant to type? If so, what does 'negation' mean to you? Hint: what are the first 3 letters of that word and what do they remind you of?

OpenStudy (anonymous):

negtive so is it x-y not equal to 10

OpenStudy (anonymous):

No.. See the example: If your age is 21 years then you are mature.. If your age is not 21 years then you are not mature..

OpenStudy (anonymous):

If the sum of x and y is 10.. then what???

OpenStudy (anonymous):

x+y=10

OpenStudy (anonymous):

Not quite. we're just trying to make the statement not true. We only need to change one thing to negate this statement.

OpenStudy (anonymous):

okay x-y=10 or x+y n ot equal 10

OpenStudy (anonymous):

not

OpenStudy (anonymous):

Right! \[x+y \neq 10\]

OpenStudy (anonymous):

so you alway change that

OpenStudy (anonymous):

can you help me with one more question/

OpenStudy (anonymous):

It really kind of depends on the problem.

OpenStudy (anonymous):

okay one min.

OpenStudy (anonymous):

what type of triangle is ABC?

OpenStudy (anonymous):

Oh my last reply was to your question of "do you always change that". I recommend closing this question and then starting a new one for your new question. :D

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yes i asked that